A 2.00-g charged particle is released from rest in a region that has a uniform electric field E= 160 N/C. After traveling a distance of 20.0 m in this region, the particle has a kinetic energy of 233 J. Determine the charge of the particle in units of coulomb.
Answer #1
2 g = .002 kg
V = E*d
V = 160*20 = 3200 V
V = EPE / q     - - - [EPE = ElectricalPotential Energy = KE ]
V*q = KE
q = KE / V = 233 / 3200 = .072813 C